A cubical object of sides 5 cm is immersed completely in water. What is the force of buoyancy on the object? (Assume g = 10 m/s2)
1
Expert's answer
2012-11-27T08:57:09-0500
Force of buoyance is F=ro*g*V
r0=1000
g=10
V=5*5*5=25*5=125*10^-6
So, F=1250000*10^-6 N=1.25N